OpenStudy (anonymous):

Using induction to prove: for n is positive integer that 8^n - 14n = 27 is divisible by 7

OpenStudy (anonymous):

so for my first step I have proven that this DOES hold true from n=1 now assume this holds true from some value of n=k. so 8^k-14k=27 is divisible by 7 show that for k+1 = 8^(k+1)- 14(k+1) +27 is divisible by 7 so 8^(k+1)= 8*8^k - 14k +27-14 i was thinking of breaking this into 8*8^k= (7+1)8^k since I know 8^k -14k+27 is divisible by 7 and 7(whatever is left) is divisible by seven.... it must be divisible by seven..... but when i try to figure it out the numbers dont come out right....

OpenStudy (zarkon):

\[8^{k+1}-14(k+1)+27\] \[8\cdot8^{k}-14k-14+27\] \[8\cdot8^{k}-(8-7)14k+(8-7)27-14\] \[8\cdot8^{k}-8\cdot 14k+8\cdot 27+7\cdot 14k-7\cdot 27-14\] \[8[8^k-14k+27]+7\cdot 14k-7\cdot 27-14\] \[8[8^k-14k+27]+7[ 14k-27-2]\]

OpenStudy (zarkon):

even better...nice work